Grade 12th passPhysical Chemistry

Wavelength=MC/h is it correct? It is false equation it correct equation

John
8 Years agoGrade 12th pass

2 Answers

Vikas Kumar
8 Years ago
E=hv
So , E=hc/wavelength; as v =c/ wavelength
Wavelength=c/v ; v=frequency and c=speed of light in vaccumn
        
Khimraj
8 Years ago
No it is not correct.
Energy E =mc2 = hc/\lambda
So \lambda = h/mc
Hope it clears...................................................